# On the Jacobs-de Leeuw-Glicksberg decomposition

### v1

@@ Abstract
For any JdLG-admissible representation $\pi$ of a semigroup $S$, we show that the reversible part is weakly equivalent to a unitary representation on a Hilbert space that decomposes into a direct sum of finite dimensional representations, and we give an alternative characterization of the almost weakly stable part in terms of the unique invariant mean on the space of weakly almost periodic functions. In the case that $S$ is a bi-amenable measured semigroup, we characterize the almost weakly stable part using invariant means and averages along Følner sequences. Moreover, we give a description of the unique projection onto the reversible part whose kernel is the almost weakly stable part in terms of ultrafilters.
@@ 1 Introduction
$\mathcal{H}_{w}$ $=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\xi\rangle|=0\right\}=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\eta\rangle|=0\ \forall\ \eta\in\mathcal{H}\right\}.$
Koopman and von Neumann [33] were the first to show that a dynamical system is weakly mixing if and only if it has no nontrivial eigenfunctions. A version of the compact-weak mixing decomposition for locally compact groups was found by Godement [27] (see also [40, Appendix 10.D]), and a version for amenable semigroups was found by Dye [18]. In between Godement and Dye, the compact-weak mixing decomposition was extended to various semigroups of operators on a Banach space by Jacobs, de Leeuw, and Glicksberg [30, 17], and is now known as the Jacobs-de Leeuw-Glicksberg (JdLG) decompoisition. Several applications of the JdLG-decomposition can be found in the literature, e.g. it is used in ergodic Ramsey theory [7, Theorem 2.3], the proof of the Erdös sumset conjecture[36], the Fürstenberg-Zimmer structure theorem [19], a proof of the Wiener-Wintner theorem [1, Proposition 2.2], and more generally in the study of weighted ergodic theorems [24, 4]. Further uses include results related to the entangled ergodic theorem [23], the Perron-Frobenius theory for completely positive operators on $W^{*}$-algebras [5], and evolution equations [25, Section V.2].
A modern treatment of the JdLG decomposition is given in [21, Chapter 16]. In the present paper we will have a heavy focus on representations of semigroups, so we give a statement of the JdLG decomposition as follows.
Theorem 1.1 (cf. [13, Theorem 6.2.18] or [28, Theorem 3.8.5]).
@@
$E_{r}$ $=\{\xi\in E\ |\ \forall\ u\in\mathscr{S},\exists\ v\in\mathscr{S}\text{ with }vu\xi=\xi\}\text{ and}$
$E_{\text{aws}}$ $=\{\xi\in E\ |\ u\xi=0\text{ for some }u\in\mathscr{S}\},$
@@
We call a representation $\pi$ of a semigroup $S$ on a Banach space $E$ JdLG-admissible if $\mathscr{T}:=(\pi_{s})_{s\in S}$ is JdLG-admissible. It is known that a relatively weakly compact semigroup $\mathscr{T}$ of operators is JdLG-admissible if and only if there exists a bi-invariant mean $M$ on $C_{b}(\mathscr{T})$ (cf. [17, Theorem 4.11]). If $S$ is a locally compact group, an amenable group, or a bi-amenable measured semigroup, then it has a unique invariant mean on $W(G)$. Consequently, the following result follows immediately from Theorem 1.1.
@@
Let $S$ be either a locally compact group, an amenable group, or a bi-amenable measured semigroup. If is $E$ a Banach space and $\pi$ is a relatively weakly compact representation of $S$ on $E$, then $\pi$ is JdLG-admissible. In particular, if $\mathscr{T}$ is a semigroup of bounded linear operators on $E$ with $c\ell_{\sigma}(\mathscr{T})=c\ell_{\sigma}(\pi(S))$, then $\mathscr{T}$ is JdLG-admissible.
@@
The main goal of the present paper is to give alternative characterizations of the reversible and almost weakly stable parts of the JdLG decomposition. Results relating the compact-weak mixing decomposition to the algebra of ultrafilters on $\mathbb{N}$ have been investigated in [8, 9, 12, 35], and in Section 5 we show that many of these results extend to JdLG-admissible semigroups of operators on a Banach space. Alternative characterizations of $E_{r}$ are already given in Proposition 16.29 and Theorem 16.31 of [21], and we will build upon these characterizations in Section 3. Alternative characterizations of $E_{\text{aws}}$ in terms of flight functions were already alluded to in [17, Corollary 5.9], [18, Theorem 1], and [28, Page 88], but were not stated explicitly prior to this paper. In particular, the next result follows from Theorem 3.4 and Lemma 4.1.
@@
    $E_{\text{aws}}=\left\{\xi\in E\ |\ \forall\ f^{\prime}\in E^{\prime},M\left(t\mapsto\left|f^{\prime}(\pi_{t}\xi)\right|\right)=0\right\}.$ (1)
When $S=\mathbb{N}$ in Theorem 1.1, alternative characterizations of $E_{\text{aws}}$ resembling the characterizations of $\mathcal{H}_{w}$ are stated in [21, Theorem 16.34]. In Theorem 4.2 we extend these alternative characterizations to various classes of amenable semigroups. For now, we only state a particularly nice corollary of Corollary 1.3, Corollary 3.2, and Theorem 4.2.
@@
$E_{aws}$ $=\left\{\xi\in E:\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{g}\xi,x^{\prime}\right\rangle\right|d\lambda(g)=0\text{ for all }x^{\prime}\in E^{\prime}\right\}$
Remark.
@@
Acknowledgements: The second author acknowledges being supported by grant 2019/34/E/ST1/00082 for the project “Set theoretic methods in dynamics and number theory,” NCN (The National Science Centre of Poland), and the first author was supported by the same grant for a 2 research visits. The second author would like to thank Uta Freiberg for funding a research visit to TU Chemnitz. We would like to thank Yulia Kuznetsova for providing us with Example 4.3. We would also like to thank Rigoberto Zelada for helpful discussions about Theorem 5.2.
@@ 2.1 Topological semigroups and weak almost periodicity
A semigroup $(S,\cdot)$ is a set $S$ with an associative binary operatation $\cdot:S\times S\rightarrow S$. If $S$ possesses an identity element $e$, then $S$ is called a monoid. If $S$ is also a Hausdorff topological space and the map $\cdot$ is continuous, then $S$ is a topological semigroup [^4] We warn the reader that in [17], the term topological semigroup refers to what we call a semitopological semigroup.. If $S$ is a Hausdorff topological space for which $\cdot$ is separately continuous, i.e., for each $s\in S$ the maps $L_{s}(t):=s\cdot t$ and $R_{s}(t):=t\cdot s$ are continuous, then $S$ is a semitopological semigroup. For the rest of this paper all semigroups are assumed to be semitopological unless it is explicitly stated otherwise. By abuse of notation, we let $R$ act on $C_{b}(S)$ and $L^{\infty}(S,\lambda)$ for some Borel measure $\lambda$, by $R_{s}(f)(t)=f(R_{s}(t))=f(t\cdot s)$, and similarly for $L$. A function $f\in C_{b}(S)$ is weakly almost periodic if $\{R_{s}f\ |\ s\in S\}$ is relatively weakly compact. We let $W(S)$ denote the space of weakly almost periodic functions on $S$. De Leeuw and Glicksberg [17, Theorem 2.7] showed that if $S$ is a compact semigroup, then $W(S)=C(S)$. For $X\in\{C_{b}(S),W(S),L^{\infty}(S,\lambda)\}$, a left-invariant mean on $X$ is a positive linear functional $m:X\rightarrow\mathbb{C}$ of norm $1$ that satisfies $m(L_{s}f)=m(f)$ for all $s\in S$ and all $f\in X$, and a right-invariant mean is defined similarly. The semigroup $S$ admits a bi-invariant mean on $X$ if there exists a left-invariant mean $m$ on $X$ that is also a right-invariant mean.
@@
In [17, Theorem 5.8] it is shown that if $W(S)$ admits a bi-invariant mean, then $W(S)$ admits a unique bi-invariant mean. It is also known that $W(S)$ admits a bi-invariant mean if and only it it admits a left-invariant mean and a (potentially different) right-invariant mean. Now let $S$ be a semigroup with bi-invariant mean $m$ on $W(S)$. If $\mathscr{S}$ is a compact semigroup and $\pi:S\rightarrow\mathscr{S}$ is a continuous homomorphism with dense image, then $\tilde{\pi}:C(\mathscr{S})\rightarrow W(S)$ is given by $(\tilde{\pi}f)(s)=f(\pi_{s})$ is well defined and $M(f):=m(\tilde{\pi}f)$ is the bi-invariant mean on $W(\mathscr{S})$ (see [17, Lemma 5.2]). There exists a compact semigroup $S^{w}$, the weakly almost periodic compactification of $S$, and a continuous injective homomorphism $i:S\rightarrow S^{w}$ with dense image, such that for any $f\in W(S)$ there is a unique $f^{w}\in W(S^{w})$ satisfying $f^{w}(i(s))=f(s)$ for all $s\in S$ (cf. [17, Theorem 5.3]). By abuse of notation, we let $i:W(S)\rightarrow W(S^{w})$ denote the map $f\mapsto f^{w}$, and we observe that $i$ is an algebra isomorphism. It is known that $W(S^{w})$ admits a bi-invariant mean $M$ if and only if $K(S^{w})$, the smallest two-sided ideal of $S^{w}$, is a compact topological group and $M(f)=\int_{K(S^{w})}fd\mu$, where $\mu$ is the normalized Haar measure on $K(S^{w})$.
@@
A group $(G,\cdot)$ is a topological group if it is a topological semigroup and the inversion map is a continuous map. All groups that we consider in this paper will be topological groups. De Leeuw and Glicksberg [17, Theorem 5.14] investigated which topological groups $G$ admit a bi-invariant mean on $W(G)$. 66 years later, Ryll-Nardzewski [39] proved a fixed point theorem that allows one to deduce that all locally compact groups $G$ admit a bi-invariant mean on $W(G)$. For further discussions on this topic the reader is referred to [28, 13, 14].
@@ 2.2 JdLG-admissible semigroups
For a Banach space $E$ we let $\mathcal{L}(E)$ denote the space of bounded linear operators on $E$. $\mathcal{L}(E)$ is naturally a semitopological semigroup when endowed with the operation of composition as well as the weak operator topology. In [21, Chapter 16.3], a semigroup $\mathscr{T}\subseteq\mathcal{L}(E)$ is called JdLG-admissible if it is relatively compact and $\mathscr{S}:=\operatorname{cl}_{\sigma}(\mathscr{T})$ contains a unique minimal idempotent $Q$. We will now show that this definition is equivalent to our definition. To this end, let us temporarily refer to the definition of JdLG-admissible from Section 1 as JdLG-decomposable in order to avoid confusion. If $\mathscr{T}$ is a relatively weakly compact semigroup, then [21, Theorem 16.5] says that the following are equivalent:
@@
[21, Theorem 16.3] tells us that every minimal right ideal of $\mathscr{S}$ is closed, and [21, Lemma 16.4] tells us that $K(\mathscr{S})$ is the union of all minimal right ideals of $\mathscr{S}$. It follows that when $\mathscr{T}$ is JdLG-admissible, $K(\mathscr{S})$ is a closed (hence compact) subsemigroup of $\mathscr{S}$. [17, Theorem 2.1] says that a compact semitopological semigroup that is algebraically a group must be a topological group. It follows that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a compact topological group. [17, Theorem 4.11] states that $K(\mathscr{S})$ is a compact topological group if and only if $\mathscr{T}$ is JdLG-decomposable. Putting everything together, we see that $\mathscr{T}$ is JdLG-decomposable if and only if it is JdLG-admissible.
@@ 2.3 Amenable semigroups and Følner nets
$\mathcal{F}$ is a two-sided Følner net if it is sumltaneously a left-Følner net and a right-Følner net. If $S$ possesses a left/right/two-sided Følner net $\mathcal{F}=(F_{i})_{i\in I}$ then $S$ is left/right/bi amenable. To see this, we define a net of functionals $(\mu_{i})_{i\in I}$ on $L^{\infty}(S,\lambda)^{\prime}$ by $\mu_{i}(f)=\frac{1}{\lambda(F_{i})}\int_{F_{i}}f(s)d\lambda(s)$ for all $f\in L^{\infty}(S,\lambda)$. Since the unit ball of $L^{\infty}(S,\lambda)^{\prime}$ is weak compact, we let $\mu$ be any cluster point of $(\mu_{i})_{i=1}^{\infty}$ and we see that $\mu$ is a left/right/two-sided invariant mean on $L^{\infty}(S,\lambda)$. We say that $\mu$ is a mean generated by $\mathcal{F}$ and we denote the set of such means by $\mathcal{M}(\mathcal{F})$.
There are examples of discrete semigroups that are amenable but do not possess a Følner net. However, if $S$ is a discrete cancellative semigroup, then $S$ is amenable if and only if it possesses a Følner net. It is also well known that if $S$ is a discrete abelian semigroup, then $S$ possesses a two-sided Følner net. We refer the reader to [37, Chapter 4.22] for further discussion on Følner conditions and amenability for discrete semigroups.
@@
If $S$ is a discrete semigroup and $\lambda$ is the counting measure, then [26, Theorem 3.15] tells us that $\overline{d}=d^{*}$ and $\underline{d}=d_{*}$. More generally, we see that $\overline{d}\leq d^{*}$ (hence $\underline{d}\geq d_{*}$) since any Følner sequences generates many means, but it is not clear if $d^{*}\leq\overline{d}$.
@@
$D(\{x\in X:f(x)\in U\})=1.$
Given a topological space $Y$, a function $f:X\rightarrow Y$, and a set $A\subseteq X$, we write
$\lim_{\underset{x\in A}{x\to\infty}}f(x)=\alpha$
if for any open neighborhood $U$ of $\alpha$ the set $\{x\in A|f(x)\notin U\}$ is compact.
A group $G$ is amenable if there exists a left-invariant mean $m_{\ell}$ on $C_{r}(G)$, the space of bounded right uniformly continuous functions on $G$. If $G$ is amenable, then there exists a right-invariant mean $m_{r}$ on $C_{\ell}(G)$ given by $m_{r}(f(g))=m_{\ell}(f(g^{-1}))$, so we do not distinguish between left and right amenability for groups. If $G$ is locally compact and $\lambda$ is a Haar measure, then $G$ is amenable if and only if $(G,\lambda)$ is amenable. Furthermore, $(G,\lambda)$ is amenable if and only if $(G,\lambda)$ possesses a left-Følner net. In any topological group $G$ we have that $W(G)\subseteq C_{r}(G)\cap C_{\ell}(G)$. It follows that if $G$ is amenable, then $W(G)$ admits a left-invariant mean and a right-invariant mean, hence it admits a bi-invariant mean. However, if $G$ is not a locally compact amenable group and $m$ is a left-invariant mean on $C_{r}(G)$, then $m$ cannot in general be extended to a left-invariant mean on $C_{b}(G)$, let alone the space of bounded Borel measurable functions on $G$. For this reasons, Theorems 4.2 and 5.2 do not apply to general amenable groups.
A group $G$ is extremely amenable if every continuous action of $G$ on a compact Hausdorff space has a fixed point. If $G$ is extremely amenable, then there exists a left invariant mean $m$ on $C_{r}(G)$ that is multiplicative in the sense that $m(fg)=m(f)m(g)$ for all $f,g\in C_{r}(G)$, which justifies the name of this class of groups. The only locally compact group that is extremely amenable is the trivial group. In particular, the class of extremely amenable groups is a naturally occurring class of amenable groups that cannot be studied with Følner nets. For further discussion on extremely amenable groups the reader is referred to [38].

### v2

@@ Abstract
For any JdLG-admissible representation $\pi$ of a semigroup $S$ on a Banach space $E$, we show that the reversible part is weakly equivalent to a unitary representation on a Hilbert space that decomposes into a direct sum of finite dimensional representations, and we give an alternative characterization of the almost weakly stable part in terms of the unique invariant mean on the space of weakly almost periodic functions. In the case that $S$ is a bi-amenable measured semigroup, we characterize the almost weakly stable part using invariant means and averages along Følner sequences. Moreover, we give a description of the unique projection onto the reversible part whose kernel is the almost weakly stable part in terms of ultrafilters.
@@ 1 Introduction
$\mathcal{H}_{w}$ $=\left\{\xi\in\mathcal{H}\ :\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\xi\rangle|=0\right\}=\left\{\xi\in\mathcal{H}\ :\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\eta\rangle|=0\ \forall\ \eta\in\mathcal{H}\right\}.$
Koopman and von Neumann [31] were the first to show that a dynamical system is weakly mixing if and only if it has no nontrivial eigenfunctions. A version of the compact-weak mixing decomposition for locally compact groups was found by Godement [26] (see also [38, Appendix 10.D]), and a version for amenable semigroups was found by Dye [17]. In between Godement and Dye, the compact-weak mixing decomposition was extended to various semigroups of operators on a Banach space by Jacobs, de Leeuw, and Glicksberg [28, 16], and is now known as the Jacobs-de Leeuw-Glicksberg (JdLG) decomposition. Several applications of the JdLG-decomposition can be found in the literature, e.g. it is used in ergodic Ramsey theory [7, Theorem 2.3], the first proof of the Erdös sumset conjecture [35], the Fürstenberg-Zimmer structure theorem [18], a proof of the Wiener-Wintner theorem [1, Proposition 2.2], and more generally in the study of weighted ergodic theorems [23, 4]. Further uses include results related to the entangled ergodic theorem [22], the Perron-Frobenius theory for completely positive operators on $W^{*}$-algebras [5], and evolution equations [24, Section V.2].
A modern treatment of the JdLG-decomposition is given in [20, Chapter 16]. In the present paper we will have a heavy focus on representations of semigroups, so we give a statement of the JdLG decomposition as follows.
Theorem 1.1 (cf. [12, Theorem 6.2.18]).
@@
$E_{r}$ $=\{\xi\in E\ :\ \forall\ u\in\mathscr{S},\exists\ v\in\mathscr{S}\text{ with }vu\xi=\xi\}\text{ and}$ (1)
$E_{\text{aws}}$ $=\{\xi\in E\ :\ u\xi=0\text{ for some }u\in\mathscr{S}\},$
@@
Remark.
We call a representation $\pi$ of a semigroup $S$ on a Banach space $E$ JdLG-admissible if $\mathscr{T}:=(\pi_{s})_{s\in S}$ is JdLG-admissible. It is known that a relatively weakly compact semigroup $\mathscr{T}$ of operators is JdLG-admissible if and only if there exists a bi-invariant mean $M$ on $C_{b}(\mathscr{T})$ (cf. [16, Theorem 4.11]). If $S$ is a topological group, or a bi-amenable measured semigroup, then it has a unique invariant mean on $W(G)$. Consequently, the next result follows immediately from Theorem 1.1.
@@
Let $S$ be either a topological group or a bi-amenable measured semigroup. If $E$ is a Banach space and $\pi$ is a relatively weakly compact representation of $S$ on $E$, then $\pi$ is JdLG-admissible. In particular, if $\mathscr{T}$ is a semigroup of bounded linear operators on $E$ with $c\ell_{\sigma}(\mathscr{T})=c\ell_{\sigma}(\pi(S))$, then $\mathscr{T}$ is JdLG-admissible.
@@
The main goal of the present paper is to give alternative characterizations of the reversible and almost weakly stable parts of the JdLG decomposition. Results relating the compact-weak mixing decomposition to the algebra of ultrafilters on $\mathbb{N}$ have been investigated in [8, 9, 11, 34], and in Section 5 we show that many of these results extend to JdLG-admissible semigroups of operators on a Banach space. Alternative characterizations of $E_{r}$ are already given in Proposition 16.29 and Theorem 16.31 of [20], and we will build upon these characterizations in Section 3. Alternative characterizations of $E_{\text{aws}}$ in terms of flight functions were already alluded to in [16, Corollary 5.9] and [17, Theorem 1], but were not stated explicitly prior to this paper. It is now known (cf. [12, Theorem 4.3.13]) that a semigroup $S$ admits a bi-invariant mean $m$ on $W(S)$ if and only if we have $W(S)=\mathcal{SAP}(S)\oplus W_{0}(S)$, where $\mathcal{SAP}(S)$ is the space of strongly almost periodic functions on $S$, and $W_{0}(S)$ consists of those functions $f\in W(S)$ that have 0 in the weak closure of the orbit of their right translates, and in this case we also have that $W_{0}(S)=\{f\in W(S)\ :\ m(|f|)=0\}$. In particular, the next result follows from Theorem 3.6 and Lemma 4.1 and can be seen as an analogue of this classical decomposition.
@@
    $E_{\text{aws}}=\left\{\xi\in E\ :\ \forall\ f^{\prime}\in E^{\prime},M\left(t\mapsto\left|f^{\prime}(\pi_{t}\xi)\right|\right)=0\right\}.$ (2)
When $S=\mathbb{N}$ in Theorem 1.1, alternative characterizations of $E_{\text{aws}}$ resembling the characterizations of $\mathcal{H}_{w}$ are stated in [20, Theorem 16.34]. In Theorem 4.2 we extend these alternative characterizations to various classes of amenable semigroups. For now, we only state a particularly nice result following from Corollary 1.3, Corollary 3.2, and Theorem 4.2.
@@
$E_{\text{aws}}$ $=\left\{\xi\in E:\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{g}\xi,x^{\prime}\right\rangle\right|d\lambda(g)=0\text{ for all }x^{\prime}\in E^{\prime}\right\}.$
@@
Acknowledgements: The second author acknowledges being supported by grant 2019/34/E/ST1/00082 for the project “Set theoretic methods in dynamics and number theory,” NCN (The National Science Centre of Poland), and the first author was supported by the same grant for a 2 research visits. The second author would like to thank Uta Freiberg for funding a research visit to TU Chemnitz. We would also like to thank Rigoberto Zelada for helpful discussions about Theorem 5.2.
@@ 2.1 Topological semigroups and weak almost periodicity
All of the facts that we collect in this section about weakly almost periodic functions come from the classical paper of de Leeuw and Glicksberg [16]. For futher discussions on this topic the reader is referred to [12, 13].
@@
A semigroup $(S,\cdot)$ is a set $S$ with an associative binary operatation $\cdot:S\times S\rightarrow S$. If $S$ possesses an identity element $e$, then $S$ is called a monoid. If $S$ is also a Hausdorff topological space and the map $\cdot$ is continuous, then $S$ is a topological semigroup [^5] We warn the reader that in [16] and [13], the term topological semigroup refers to what we call a semitopological semigroup.. If $S$ is a Hausdorff topological space for which $\cdot$ is separately continuous, i.e., for each $s\in S$ the maps $L_{s}(t):=s\cdot t$ and $R_{s}(t):=t\cdot s$ are continuous, then $S$ is a semitopological semigroup. For the rest of this paper all semigroups are assumed to be semitopological unless it is explicitly stated otherwise. By abuse of notation, we let $R$ act on $C_{b}(S)$ and on $L^{\infty}(S,\lambda)$ for some Borel measure $\lambda$, by $R_{s}(f)(t)=f(R_{s}(t))=f(t\cdot s)$, and similarly for $L$. A function $f\in C_{b}(S)$ is weakly almost periodic if $\{R_{s}f\ :\ s\in S\}$ is relatively weakly compact. We let $W(S)$ denote the space of weakly almost periodic functions on $S$. If $S$ is a compact semigroup, then $W(S)=C(S)$. For $X\in\{C_{b}(S),W(S),L^{\infty}(S,\lambda)\}$, a left-invariant mean on $X$ is a positive linear functional $m:X\rightarrow\mathbb{C}$ of norm $1$ that satisfies $m(L_{s}f)=m(f)$ for all $s\in S$ and all $f\in X$, and a right-invariant mean is defined similarly. The semigroup $S$ admits a bi-invariant mean on $X$ if there exists a left-invariant mean $m$ on $X$ that is also a right-invariant mean.
@@
If $W(S)$ admits a bi-invariant mean, then it is unique. It is also known that $W(S)$ admits a bi-invariant mean if and only if it admits a left-invariant mean and a (potentially different) right-invariant mean. Now let $S$ be a semigroup with a bi-invariant mean $m$ on $W(S)$. If $\mathscr{S}$ is a compact semigroup and $\pi:S\rightarrow\mathscr{S}$ is a continuous homomorphism with dense image, then $\tilde{\pi}:C(\mathscr{S})\rightarrow W(S)$ is given by $(\tilde{\pi}f)(s)=f(\pi_{s})$ is well defined and $M(f):=m(\tilde{\pi}f)$ is the bi-invariant mean on $W(\mathscr{S})=C(\mathscr{S})$. There exists a compact semigroup $S^{w}$, the weakly almost periodic compactification of $S$, and a continuous injective homomorphism $i:S\rightarrow S^{w}$ with dense image, such that for any $f\in W(S)$ there is a unique $f^{w}\in W(S^{w})$ satisfying $f^{w}(i(s))=f(s)$ for all $s\in S$. By abuse of notation, we let $i:W(S)\rightarrow W(S^{w})$ denote the map $f\mapsto f^{w}$, and we observe that $i$ is an algebra isomorphism. It is known that $W(S^{w})$ admits a bi-invariant mean $M$ if and only if $K(S^{w})$, the smallest two-sided ideal of $S^{w}$, is a compact topological group and $M(f)=\int_{K(S^{w})}fd\mu$, where $\mu$ is the normalized Haar measure on $K(S^{w})$.
@@
A group $(G,\cdot)$ is a topological group if it is a topological semigroup and the inversion map is a continuous map. All groups that we consider in this paper will be topological groups. De Leeuw and Glicksberg [16] investigated which topological groups $G$ admit a bi-invariant mean on $W(G)$, and $6$ years later, Ryll-Nardzewski [37] proved a fixed point theorem that allows one to deduce that all topological groups $G$ admit a bi-invariant mean on $W(G)$.
@@ 2.2 JdLG-admissible semigroups
For a Banach space $E$ we let $\mathcal{L}(E)$ denote the space of bounded linear operators on $E$. $\mathcal{L}(E)$ is naturally a semitopological semigroup when endowed with the operation of composition as well as the weak operator topology. In [20, Chapter 16.3], a semigroup $\mathscr{T}\subseteq\mathcal{L}(E)$ is called JdLG-admissible if it is relatively compact and $\mathscr{S}:=\operatorname{cl}_{\sigma}(\mathscr{T})$ contains a unique minimal idempotent $Q$. We will now show that this definition is equivalent to our definition. To this end, let us temporarily refer to the definition of JdLG-admissible from Section 1 as JdLG-decomposable in order to avoid confusion. If $\mathscr{T}$ is a relatively weakly compact semigroup, then [20, Theorem 16.5] says that the following are equivalent:
@@
[20, Theorem 16.3] tells us that every minimal right ideal of $\mathscr{S}$ is closed, and [20, Lemma 16.4] tells us that $K(\mathscr{S})$ is the union of all minimal right ideals of $\mathscr{S}$. It follows that when $\mathscr{T}$ is JdLG-admissible, $K(\mathscr{S})$ is a closed (hence compact) subsemigroup of $\mathscr{S}$. [16, Theorem 2.1] says that a compact semitopological semigroup that is algebraically a group must be a topological group. It follows that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a compact topological group. [16, Theorem 4.11] states that $K(\mathscr{S})$ is a compact topological group if and only if $\mathscr{T}$ is JdLG-decomposable. Putting everything together, we see that $\mathscr{T}$ is JdLG-decomposable if and only if it is JdLG-admissible.
@@ 2.3 Amenable semigroups and Følner nets
$\mathcal{F}$ is a two-sided Følner net if it is simultaneously a left-Følner net and a right-Følner net. If $S$ possesses a left/right/two-sided Følner net $\mathcal{F}=(F_{i})_{i\in I}$ then $S$ is left/right/bi amenable. To see this, we define a net of functionals $(\mu_{i})_{i\in I}$ on $L^{\infty}(S,\lambda)^{\prime}$ by $\mu_{i}(f)=\frac{1}{\lambda(F_{i})}\int_{F_{i}}f(s)d\lambda(s)$ for all $f\in L^{\infty}(S,\lambda)$. Since the unit ball of $L^{\infty}(S,\lambda)^{\prime}$ is weak compact, we let $\mu$ be any cluster point of $(\mu_{i})_{i=1}^{\infty}$ and we see that $\mu$ is a left/right/two-sided invariant mean on $L^{\infty}(S,\lambda)$. We say that $\mu$ is a mean generated by $\mathcal{F}$ and we denote the set of such means by $\mathcal{M}(\mathcal{F})$.
There are examples of discrete semigroups that are amenable but do not possess a Følner net. However, if $S$ is a discrete cancellative semigroup, then $S$ is amenable if and only if it possesses a Følner net. It is also well known that if $S$ is a discrete abelian semigroup, then $S$ possesses a two-sided Følner net. We refer the reader to [36, Chapter 4.22] for further discussion on Følner conditions and amenability for discrete semigroups.
@@
If $S$ is a discrete semigroup and $\lambda$ is the counting measure, then [25, Theorem 3.15] tells us that $\overline{d}=d^{*}$ and $\underline{d}=d_{*}$. More generally, we see that $\overline{d}\leq d^{*}$ (hence $\underline{d}\geq d_{*}$) since any Følner sequences generates many means, but it is not clear if $d^{*}\leq\overline{d}$.
@@
$D(\{s\in S\ :\ f(s)\in U\})=1.$
Given a topological space $Y$, a function $f:S\rightarrow Y$, and a set $A\subseteq S$, we write
$\lim_{\underset{s\in A}{s\to\infty}}f(s)=\alpha$
if for any open neighborhood $U$ of $\alpha$ the set $\{s\in A\ :\ f(s)\notin U\}$ is contained in a compact set.